# ViSTA: Visual Storytelling using Multi-modal Adapters for Text-to-Image Diffusion Models

### v2

@@ Appendix B Limitations
Figure 5. ViSTA sample on FlintStonesSV test set.
Figure 6. ViSTA sample on FlintStonesSV test set.
@@
Figure 7. Additional Qualitative results. This figure presents a comparison of storytelling between ViSTA, baseline methods, and state-of-the-art on a sample StorySalon story. While SDXL-Prompt show high-quality and well-aligned images, they fail in generating consistent character acorss all frames. Although IP-Adapter shows consistent character, the generated images do not align with the prompt. Compare with the state-of-the-art StoryGen, our ViSTA shows better consistency on both characters and style.
@@
Figure 8. Additional Qualitative results. This figure presents a comparison of storytelling between ViSTA, baseline methods, and state-of-the-art on a sample StorySalon story. While SDXL-Prompt show high-quality and well-aligned images, they fail in generating consistent character acorss all frames. Although IP-Adapter shows consistent character, the generated images do not align with the prompt. Compare with the state-of-the-art StoryGen, our ViSTA shows better consistency on both characters and style.
@@
Figure 9. Additional Qualitative results. This figure presents a comparison of storytelling between ViSTA, baseline methods, and state-of-the-art on a sample StorySalon story. While SDXL-Prompt show high-quality and well-aligned images, they fail in generating consistent character acorss all frames. Although IP-Adapter shows consistent character, the generated images do not align with the prompt. Compare with the state-of-the-art StoryGen, our ViSTA shows better consistency on both characters and style.
@@
Figure 10. Additional Qualitative results. This figure presents a comparison of storytelling between ViSTA, baseline methods, and state-of-the-art on a sample StorySalon story. While SDXL-Prompt show high-quality and well-aligned images, they fail in generating consistent character acorss all frames. Although IP-Adapter shows consistent character, the generated images do not align with the prompt. Compare with the state-of-the-art StoryGen, our ViSTA shows better consistency on both characters and style.

### v3

@@ Appendix D Limitations
Figure 9: Additional Qualitative results. This figure presents a comparison of storytelling between ViSTA, baseline methods, and state-of-the-art on a sample StorySalon story. While SDXL-Prompt show high-quality and well-aligned images, they fail in generating consistent character acorss all frames. Although IP-Adapter shows consistent character, the generated images do not align with the prompt. Compare with the state-of-the-art StoryGen, our ViSTA shows better consistency on both characters and style.
@@
Figure 10: Additional Qualitative results. This figure presents a comparison of storytelling between ViSTA, baseline methods, and state-of-the-art on a sample StorySalon story. While SDXL-Prompt show high-quality and well-aligned images, they fail in generating consistent character acorss all frames. Although IP-Adapter shows consistent character, the generated images do not align with the prompt. Compare with the state-of-the-art StoryGen, our ViSTA shows better consistency on both characters and style.
@@
Figure 11: Additional Qualitative results. This figure presents a comparison of storytelling between ViSTA, baseline methods, and state-of-the-art on a sample StorySalon story. While SDXL-Prompt show high-quality and well-aligned images, they fail in generating consistent character acorss all frames. Although IP-Adapter shows consistent character, the generated images do not align with the prompt. Compare with the state-of-the-art StoryGen, our ViSTA shows better consistency on both characters and style.
@@
Figure 12: Additional Qualitative results. This figure presents a comparison of storytelling between ViSTA, baseline methods, and state-of-the-art on a sample StorySalon story. While SDXL-Prompt show high-quality and well-aligned images, they fail in generating consistent character acorss all frames. Although IP-Adapter shows consistent character, the generated images do not align with the prompt. Compare with the state-of-the-art StoryGen, our ViSTA shows better consistency on both characters and style.